# Maxitive functions with respect to general orders

### v2

@@ 2. Basic concepts and definitions
In general, for an arbitrary $\mathcal{Y}$-directed subset $A\subset X$, it does not follow that $\alpha_{\min}(\phi):=\sup_{x\in A}\phi(x)$ is $\mathcal{Y}$-stable, even if $X$ is $\mathcal{Y}$-complete.
To give a counter example, let $\mathcal{Y}$ be the family of all at most countable subsets, and consider the set $X$ of all functions $f\colon\mathbb{R}\to[0,1]$ such that there exists a bounded interval $I\subset\mathbb{R}$ for which $\{x\in\mathbb{R}\setminus I\colon f(x)>0\}$ is at most countable. We consider the family $\Phi:=\{\phi_{x}\colon x\in\mathbb{R}\}$, where $\phi_{x}(f)=f(x)$. The ordered set $(X,\preccurlyeq)$ is $\mathcal{Y}$-complete. Now, let $A$ be the set of all functions $f\in X$ such that $\{x\in\mathbb{R}\colon f(x)>0\}$ is at most countable. Then the set $A$ is $\mathcal{Y}$-directed and $\alpha_{\min}(\phi_{x})=\sup_{f\in A}f(x)=1$ for all $x\in\mathbb{R}$. To show that $\alpha_{\min}$ is not $\mathcal{Y}$-stable, consider for each $n\in\mathbb{N}$ the indicator function $1_{[-n,n]}$ which assumes the value $1$ on $[-n,n]$ and 00 on $[-n,n]^{c}$. Then it holds $\sup_{n\in\mathbb{N}}1_{[-n,n]}(x)\leq\alpha_{\min}(\phi_{x})$ for all $x\in\mathbb{R}$, but there is no function $f\in X$ with $1_{[-n,n]}(x)\leq f(x)$ for all $x\in\mathbb{R}$.
@@ 3. Main result
As before, let $X$ be a non-empty set, and let $\Phi$ be a non-empty set of functions $\phi\colon X\to\mathbb{R}$ which induces the preorder $x\preccurlyeq y$ if and only if $\phi(x)\leq\phi(y)$ for all $\phi\in\Phi$. From now on, if not stated otherwise, we assume in addition that
  • the set $\Phi$ is endowed with a topology such that $\Phi$ is a Polish space, and the functions $f_{x}\colon\Phi\to\mathbb{R}$ defined by $f_{x}(\phi):=\phi(x)$ are lower semicontinuous for all $x\in X$;
  • $\mathcal{Y}$ is the family of all at most countable subsets.
In that case, we denote the properties $\mathcal{Y}$-complete, $\mathcal{Y}$-maxitive, $\mathcal{Y}$-directed and $\mathcal{Y}$-stable as countably complete, countably maxitive, countably directed and countably stable, respectively. Note that we do not assume that $(X,\preccurlyeq)$ is countably complete.
The following is our main result.
@@
Let $\psi\colon X\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$ given by $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$. Then the following statements are equivalent:
@@
    $\psi$ is monotone and has countably directed sublevel sets.
@@
Then the mapping $F\colon I\times L(\Phi)\to[-\infty,\infty]$ is well-defined and satisfies the following properties. For every $f,g\in L(\Phi)$ and $s,t\in I$, it holds
@@
To show (4), let $(f_{n})_{n\in\mathbb{N}}\subset L(\Phi)$ with $f:=\sup_{n\in\mathbb{N}}f_{n}\in L(\Phi)$. Then it follows from condition (2) that $\sup_{n\in\mathbb{N}}F(s,f_{n})\leq F(s,f)$. By contradiction, we assume there exists $c\in\mathbb{R}$ with
@@
Then for every $n\in\mathbb{N}$ there exists $x_{n}\in A_{s}$ with $f_{n}(\phi)-\phi(x_{n})<c$ for all $\phi\in\Phi$. Since $A_{s}$ is countably directed, there exists $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$. Using (2.1), we obtain $f_{n}(\phi)-\phi(x^{\ast})<c$ for all $\phi\in\Phi$ and $n\in\mathbb{N}$. Taking the supremum first over all $n\in\mathbb{N}$, and second over all $\phi\in\Phi$, we obtain
@@
To verify (5), suppose by contradiction that $F(s,f)=-\infty$. Then for $\phi\in\Phi$ there exists $(x_{n})_{n\in\mathbb{N}}\subset A_{s}$ with $f(\phi)-\phi(x_{n})\leq-n$ for all $n\in\mathbb{N}$. Since $A_{s}$ is countably directed, there is $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$, and therefore
@@
Letting $n\to\infty$ yields $\phi(x^{\ast})=\infty$ in contrast to the assumption that $\phi$ is real-valued. This shows that $F(s,f)\in(-\infty,\infty]$. In particular, we obtain for every $x\in A_{s}$,
@@
If $x\in X$ satisfies $F(s,f_{x})\leq 0$, then there exists $(x_{n})_{n\in\mathbb{N}}\subset A_{s}$ with $\phi(x)-\phi(x_{n})<1/n$ for all $\phi\in\Phi$. Since $A_{s}$ is countably directed, there exists $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$, and therefore $\phi(x)-\phi(x^{\ast})\leq 0$ for all $\phi\in\Phi$, which shows $x\preccurlyeq x^{\ast}$. Since $\psi$ is monotone, we conclude that $\psi(x)\leq\psi(x^{\ast})\leq s$, and therefore $x\in A_{s}$.
@@
where $\alpha(s,\phi):=\sup_{f\in C_{b}(\Phi)}\{f(\phi)-F(s,f)\}$ for all $s\in I$ and $\phi\in\Phi$. As a result of condition (7), we obtain
@@
where we set $\alpha(s,\phi):=-\infty$ whenever $s\notin I$. Moreover, the infimum is attained whenever $\psi(x)<\infty$. Due to condition (1), the mapping $\alpha$ is non-decreasing in the first argument. For $s\in\mathbb{R}$, the representation (3.7) implies that $x\in A_{s}$ if and only if $\phi(x)\leq\alpha(s,\phi)$ for all $\phi\in\Phi$. Since $A_{s}$ is countably directed, it follows from Proposition 2.5 that $\alpha$ is countably stable in the second argument.
$(ii)\Rightarrow(iii)$: Fix $s\in\mathbb{R}$. The representation (3.1) implies that $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha(s,\phi)\}\leq 0$. Since $\alpha$ is countably stable in the second argument, it follows from Proposition 2.5 that $A_{s}$ is countably directed. Moreover, Proposition 2.5 implies that $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$, where $\alpha_{\min}(s,\phi):=\sup_{x\in A_{s}}\phi(x)$, and $\alpha_{\min}$ is countably stable in the second argument.
$(iii)\Rightarrow(iv)$: Fix $s\in\mathbb{R}$. It follows from (iii) together with Proposition 2.5 that $A_{s}$ is countably directed and $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$. Hence, $x\in X$ satisfies $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$, if and only if $x\in A_{t}$ for all $t>s$, if and only if $\sup_{t>s}\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(t,\phi)\}\leq 0$, if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha^{+}_{\min}(t,\phi)\}\leq 0$. In particular, it follows from Proposition 2.5 that $\alpha^{+}$ is coutable stable in the second argument.
@@
This shows that $\psi(x)=\sup_{\phi\in\Phi}G(\phi(x),\phi)$ for all $x\in X$. In case that $\psi(x)=\infty$, then the representation (3.3) implies that for every $n\in\mathbb{N}$ we can find $\phi_{n}\in\Phi$ such that $n\notin\alpha_{\phi}^{-1}([\phi_{n}(x),\infty))$. Hence, for every $n\in\mathbb{N}$,
@@
$(v)\Rightarrow(i)$ Since $\alpha^{+}_{\min}$ is countably stable in the second argument, we argue similar as in the implication from (iii) to (iv) that the level sets of $\psi$ are countably directed. Since $G$ is non-decreasing in the first argument, the functional $\psi$ is monotone. ∎
@@
We point out that the statement of Theorem 3.1 remains valid if we replace countable maxitivity by complete maxitivity and countable stability by complete stability. [^3] I.e., $\mathcal{P}(X)$-maxitivity and $\mathcal{P}(X)$-stability, where $\mathcal{P}(X)$ is the power set. The proof follows with the same arguments and the following modifications. In the implication $(i)\Rightarrow(ii)$, we do not rely on Corollary A.3, but rather extend the definition (3.5) of $F(s,f)$ for all $s\in I$ and $f\in[-\infty,\infty)^{\Phi}$, where $[-\infty,\infty)^{\Phi}$ denotes the set of all functions from $\Phi$ to [,)[$\infty$,$\infty$)[ - ∞ , ∞ ). Direct verification shows that $F$ satisfies the conditions (1)–(3) on $I\times[-\infty,\infty)^{\Phi}$ and $F(s,\sup_{j\in J}f_{j})=\sup_{j\in J}F(s,f_{j})$ for every family $(f_{j})_{j\in J}$ in $[-\infty,\infty)^{\Phi}$. From the latter property, one obtains the representation (3.6) for all $s\in I$ and $f\in[-\infty,\infty)^{\Phi}$ for a suitable function $\alpha\colon I\times\Phi\to(-\infty,\infty]$. From there, the argumentation is similar as in the proof of $(i)\Rightarrow(ii)$ in Theorem 3.1.
For the rest of this section, the space $X$ is endowed with an operation $+\colon X\times\mathbb{R}\to X$. A function $\psi\colon X\to[-\infty,\infty]$ satisfies the translation property if $\psi(x+c)=\psi(x)+c$ for all $x\in X$ and $c\in\mathbb{R}$. A function $\psi\colon X\to[-\infty,\infty]$ satisfies the deviation property if $\psi(x+c)=\psi(x)$ for all $x\in X$ and $c\in\mathbb{R}$.
@@
Suppose that every $\phi\in\Phi$ has the translation property. Let $\psi\colon X\to(-\infty,\infty]$ be a function with $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$ for all $s\in\mathbb{R}$. Then the following statements are equivalent:
@@
    $\psi$ is monotone and has countably directed sublevel sets.
@@
    There exists a stable function $\gamma\colon\Phi\to[-\infty,\infty]$ with
    $\psi(x)=\sup_{\phi\in\Phi}\{\phi(x)-\gamma(\phi)\}\quad\mbox{for all }x\in X.$
@@
    The mapping $\phi\mapsto\alpha_{\min}(0,\phi):=\sup_{x\in A_{0}}\psi(x)$ is stable and
    $\psi(x)=\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(0,\phi)\}\quad\mbox{for all }x\in X.$
@@
$(i)\Rightarrow(iii)$: Since $\psi$ has the translation property, it holds $\psi(x)\leq s$ if and only if $\psi(x-s)\leq 0$, so that $A_{s}=A_{0}-s$ for all $s\in\mathbb{R}$. Hence, since every $\phi\in\Phi$ has the translation property, we obtain for every $s\in\mathbb{R}$ and $\phi\in\Phi$,
@@
$(ii)\Rightarrow(i)$: That $\psi$ is monotone and has countably directed sublevel sets follows from Theorem 3.1. Moreover, since every $\phi\in\Phi$ has the translation property, for every $x\in X$ and $c\in\mathbb{R}$,
@@
Suppose that $\psi\colon X\to(-\infty,\infty]$ is monotone and has countably directed sublevel sets. If $\psi$ has the deviation property, then $\psi(x)=\infty$ for all $x\in X$.
@@
4. Examples
Let $(\Omega,\mathcal{F},\mathbb{P})$ be an atomless probability space. We denote by $\mathcal{M}$ the set of all random variables $\xi\colon\Omega\to\mathbb{R}$, where two of them are identified if they have the same law. Further, let $\mathcal{M}_{\infty}$ and $\mathcal{M}_{1}$ be the sets of all random variables in $\mathcal{M}$ which are essentially bounded, and have finite first moments, respectively. In the following, we analyze the main results of the previous section for several stochastic orders. For definitions and further details on stochastic orders we refer to [16, 22].

### v3

@@ 3. Basic concepts and definitions
In general, for an arbitrary countably upwards directed subset $A\subset X$, it does not follow that $\alpha_{\min}(\phi):=\sup_{x\in A}\phi(x)$ is countably stable, even if $X$ is countably complete.
To give a counterexample, let $X$ be the set of all functions $f\colon\mathbb{R}\to[0,1]$ such that there exists a bounded interval $I\subset\mathbb{R}$ for which $\{x\in\mathbb{R}\setminus I\colon f(x)>0\}$ is at most countable. We consider the family $\Phi:=\{\phi_{x}\colon x\in\mathbb{R}\}$, where $\phi_{x}(f)=f(x)$. The ordered set $(X,\preccurlyeq)$ is countably complete. Now, let $A$ be the set of all functions $f\in X$ such that $\{x\in\mathbb{R}\colon f(x)>0\}$ is at most countable. Then, the set $A$ is countably upwards directed, and
$\alpha_{\min}(\phi_{x})=\sup_{f\in A}f(x)=1\quad\text{for all }x\in\mathbb{R}.$
To show that $\alpha_{\min}$ is not countably stable, consider for every $n\in\mathbb{N}$ the indicator function $1_{[-n,n]}$ which assumes the value $1$ on $[-n,n]$ and 00 on $[-n,n]^{c}$. Then, it holds
$\sup_{n\in\mathbb{N}}1_{[-n,n]}(x)\leq\alpha_{\min}(\phi_{x})\quad\text{for all }x\in\mathbb{R},$
but there is no function $f\in X$ with $1_{[-n,n]}(x)\leq f(x)$ for all $x\in\mathbb{R}$.
@@ 4. Main result
In this section, we present our representation results for maxitive functions with respect to general orders. Since we do not assume any form of completeness on the order, we consider monotone functions with countably upwards directed sublevel sets. In contrast to bounded countable maxitivity, these concepts do not require that the order is countably complete, allowing us to derive more general statements without relying on the assumption of countable completeness.
Throughout this section, we assume that $X$ is a non-empty set, and $\Phi$ is a non-empty set of functions $\phi\colon X\to\mathbb{R}$ which induces the preorder
$x\preccurlyeq y\quad\text{if and only if}\quad\phi(x)\leq\phi(y)\text{ for all}\phi\in\Phi.$
In addition, we assume that the set $\Phi$ is endowed with a topology such that $\Phi$ is a Polish space, and the functions $f_{x}\colon\Phi\to\mathbb{R}$ defined by $f_{x}(\phi):=\phi(x)$ are lower semicontinuous [^15] I.e., $\{\phi\in\Phi\colon f_{x}(\phi)\leq s\}$ is closed for all $s\in\mathbb{R}$. for all $x\in X$.
The following is the main result of this article.
@@
Let $\psi\colon X\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$ given by
$A_{s}:=\{x\in X\colon\psi(x)\leq s\}.$
Then, the following statements are equivalent:
@@
    $\psi$ is monotone and has countably upwards directed sublevel sets.
@@
Then, the mapping $F\colon I\times L(\Phi)\to[-\infty,\infty]$ is well-defined and satisfies the following properties. For every $f,g\in L(\Phi)$ and $s,t\in I$, it holds
@@
To show (4), let $(f_{n})_{n\in\mathbb{N}}\subset L(\Phi)$ with $f:=\sup_{n\in\mathbb{N}}f_{n}\in L(\Phi)$. Then, it follows from condition (2) that $\sup_{n\in\mathbb{N}}F(s,f_{n})\leq F(s,f)$. By contradiction, we assume there exists $c\in\mathbb{R}$ with
@@
Then for every $n\in\mathbb{N}$ there exists $x_{n}\in A_{s}$ with $f_{n}(\phi)-\phi(x_{n})<c$ for all $\phi\in\Phi$. Since $A_{s}$ is countably upwards directed, there exists $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$. We obtain that $f_{n}(\phi)-\phi(x^{\ast})<c$ for all $\phi\in\Phi$ and $n\in\mathbb{N}$. Taking the supremum first over all $n\in\mathbb{N}$, and second over all $\phi\in\Phi$, we obtain that
@@
To verify (5), suppose by contradiction that $F(s,f)=-\infty$. Then, for $\phi\in\Phi$ there exists $(x_{n})_{n\in\mathbb{N}}\subset A_{s}$ with $f(\phi)-\phi(x_{n})\leq-n$ for all $n\in\mathbb{N}$. Since $A_{s}$ is countably upwards directed, there is $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$, and therefore
@@
Letting $n\to\infty$ yields $\phi(x^{\ast})=\infty$ in contrast to the assumption that $\phi$ is real-valued. This shows that $F(s,f)\in(-\infty,\infty]$. In particular, we obtain that for every $x\in A_{s}$,
@@
If $x\in X$ satisfies $F(s,f_{x})\leq 0$, then there exists $(x_{n})_{n\in\mathbb{N}}\subset A_{s}$ with $\phi(x)-\phi(x_{n})<1/n$ for all $\phi\in\Phi$. Since $A_{s}$ is countably upwards directed, there exists $x^{\ast}\in A_{s}$ with $x_{1},x_{2},\cdots\preccurlyeq x^{\ast}$, and therefore $\phi(x)-\phi(x^{\ast})\leq 0$ for all $\phi\in\Phi$, which shows $x\preccurlyeq x^{\ast}$. Since $\psi$ is monotone, we conclude that $\psi(x)\leq\psi(x^{\ast})\leq s$, and therefore $x\in A_{s}$.
@@
where $\alpha(s,\phi):=\sup_{f\in C_{b}(\Phi)}\{f(\phi)-F(s,f)\}$ for all $s\in I$ and $\phi\in\Phi$. As a result of condition (7), we obtain that
@@
where we set $\alpha(s,\phi):=-\infty$ whenever $s\notin I$. Moreover, the infimum is attained whenever $\psi(x)<\infty$. Due to condition (1), the mapping $\alpha$ is non-decreasing in the first argument. For $s\in\mathbb{R}$, the representation (4.7) implies that $x\in A_{s}$ if and only if $\phi(x)\leq\alpha(s,\phi)$ for all $\phi\in\Phi$. Since $A_{s}$ is countably upwards directed, it follows from Proposition 3.6 that $\alpha$ is countably stable in the second argument.
$(ii)\Rightarrow(iii)$: Fix $s\in\mathbb{R}$. The representation (4.1) implies that $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha(s,\phi)\}\leq 0$. Since $\alpha$ is countably stable in the second argument, it follows from Proposition 3.6 that $A_{s}$ is countably upwards directed. Moreover, Proposition 3.6 implies that $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$, where $\alpha_{\min}(s,\phi):=\sup_{x\in A_{s}}\phi(x)$, and $\alpha_{\min}$ is countably stable in the second argument.
$(iii)\Rightarrow(iv)$: Fix $s\in\mathbb{R}$. It follows from (iii) together with Proposition 3.6 that $A_{s}$ is countably directed and $x\in A_{s}$ if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$. Hence, $x\in X$ satisfies $\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(s,\phi)\}\leq 0$, if and only if $x\in A_{t}$ for all $t>s$, if and only if $\sup_{t>s}\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(t,\phi)\}\leq 0$, if and only if $\sup_{\phi\in\Phi}\{\phi(x)-\alpha^{+}_{\min}(t,\phi)\}\leq 0$. In particular, it follows from Proposition 3.6 that $\alpha^{+}$ is countable stable in the second argument.
@@
This shows $\psi(x)=\sup_{\phi\in\Phi}G(\phi(x),\phi)$ for all $x\in X$. If $\psi(x)=\infty$, then the representation (4.3) implies that for every $n\in\mathbb{N}$, there is $\phi_{n}\in\Phi$ such that $n\notin\alpha_{\phi}^{-1}([\phi_{n}(x),\infty))$. Hence, for every $n\in\mathbb{N}$,
@@
$(v)\Rightarrow(i)$ Since $\alpha^{+}_{\min}$ is countably stable in the second argument, we argue as in the implication from (iii) to (iv) that the level sets of $\psi$ are countably upwrads directed. Since $G$ is non-decreasing in the first argument, the function $\psi$ is monotone. ∎
@@
A modification of the previous result also leads to the representation of bounded completely maxitive functions [^17] Defined similarly to Definition 3.3, but for arbitrary subsets rather than countable subsets.. Indeed, Theorem 4.1 remains valid if we replace countably upwards directed sublevel sets with completely upwards directed sublevel sets [^18] I.e., for every $s\in\mathbb{R}$, there exists $y^{\ast}\in A_{s}$ with $y\preccurlyeq y^{\ast}$ for all $y\in A_{s}$. and countable stability with complete stability [^19] A function $\alpha\colon{\Phi}\to[-\infty,\infty]$ is completely stable if for every subset $Y\subset X$ with $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, there exists $y^{\ast}\in X$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$ and $\phi(y^{\ast})\leq\alpha(\phi)$ for all $\phi\in\Phi$.. The proof follows with the same arguments, subject to the following modifications. In the implication $(i)\Rightarrow(ii)$, we do not rely on Corollary A.3, but rather extend the definition (4.5) of $F(s,f)$ for all $s\in I$ and $f\in[-\infty,\infty)^{\Phi}$, where $[-\infty,\infty)^{\Phi}$ denotes the set of all functions from $\Phi$ to [,)[$\infty$,$\infty$)[ - ∞ , ∞ ). Direct verification shows that $F$ satisfies the conditions (1)–(3) on $I\times[-\infty,\infty)^{\Phi}$ and $F(s,\sup_{j\in J}f_{j})=\sup_{j\in J}F(s,f_{j})$ for every family $(f_{j})_{j\in J}$ in $[-\infty,\infty)^{\Phi}$. From the latter property, one obtains the representation (4.6) for all $s\in I$ and $f\in[-\infty,\infty)^{\Phi}$ for a suitable function $\alpha\colon I\times\Phi\to(-\infty,\infty]$. From there, the argumentation is similar as in the proof of $(i)\Rightarrow(ii)$ in Theorem 4.1.
Remark 4.3.
In line with (2.1), (2.2), and (2.3), the function $\alpha$ in (4.1) can be interpreted as a penalty function. In the present setting, there may exist different penalties $\alpha$ for which the representation (4.1) holds.
We note that $\alpha_{\min}$ is the smallest penalty $\alpha$ for which the representation (4.1) is valid. Indeed, if (4.1) holds for a penalty $\alpha$, then for any fixed $\phi\in\Phi$ and $s\in\mathbb{R}$, it follows that $\phi(x)\leq\alpha(s,\phi)$ for all $x\in A_{s}$. Taking the supremum over all $x\in A_{s}$, we obtain that $\alpha_{\min}(s,\phi)=\sup_{x\in A_{s}}\phi(x)\leq\alpha(s,\phi)$. Furthermore, by construction, $\alpha_{\min}^{+}$ is the smallest penalty for which the representation (4.1) holds and is also right-continuous in the first argument. Note that the right continuity of $\alpha_{\min}^{+}$ is necessary in the implication $(iv)\Rightarrow(v)$ in Theorem 4.1.
For the remainder of this section, $X$ is endowed with an operation $+\colon X\times\mathbb{R}\to X$. A function $\psi\colon X\to[-\infty,\infty]$ satisfies the translation property if $\psi(x+c)=\psi(x)+c$ for all $x\in X$ and $c\in\mathbb{R}$. A function $\psi\colon X\to[-\infty,\infty]$ satisfies the deviation property if $\psi(x+c)=\psi(x)$ for all $x\in X$ and $c\in\mathbb{R}$.
@@
Suppose that every $\phi\in\Phi$ has the translation property. Let $\psi\colon X\to(-\infty,\infty]$ be a function with $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$ for all $s\in\mathbb{R}$. Then, the following statements are equivalent:
@@
    $\psi$ has the translation property, is monotone, and has countably upwards directed sublevel sets.
@@
    There exists a countably stable function $\gamma\colon\Phi\to[-\infty,\infty]$ with
    $\psi(x)=\sup_{\phi\in\Phi}\{\phi(x)-\gamma(\phi)\}\quad\text{for all }x\in X.$
@@
    The mapping $\phi\mapsto\alpha_{\min}(0,\phi):=\sup_{x\in A_{0}}\phi(x)$ is countably stable and
    $\psi(x)=\sup_{\phi\in\Phi}\{\phi(x)-\alpha_{\min}(0,\phi)\}\quad\text{for all }x\in X.$
@@
$(i)\Rightarrow(iii)$: Since $\psi$ has the translation property, it holds $\psi(x)\leq s$ if and only if $\psi(x-s)\leq 0$, so that $A_{s}=A_{0}-s:=\{x-s\colon x\in A_{0}\}$ for all $s\in\mathbb{R}$. Hence, since every $\phi\in\Phi$ has the translation property, we obtain for every $s\in\mathbb{R}$ and $\phi\in\Phi$,
@@
$(ii)\Rightarrow(i)$: That $\psi$ is monotone and has countably upwards directed sublevel sets follows from Theorem 4.1. Moreover, since every $\phi\in\Phi$ has the translation property, for every $x\in X$ and $c\in\mathbb{R}$,
@@
Suppose that every $\phi\in\Phi$ satisfies the deviation property, and let $\psi\colon X\to(-\infty,\infty]$ be a monotone function with countably upwards directed sublevel sets. If $\psi$ additionally satisfies the translation property, then $\psi(x)=\infty$ for all $x\in X$.
@@
5. Maxitivity with respect to stochastic orders
Let $(\Omega,\mathcal{F},\mathbb{P})$ be an atomless probability space. [^20] Recall that a probability space $(\Omega,\mathcal{F},\mathbb{P})$ is atomless if for all $A\in\mathcal{F}$ with $\mathbb{P}(A)>0$ there exists $B\in\mathcal{F}$ such that $B\subset A$ and $0<\mathbb{P}(B)<\mathbb{P}(A)$; see e.g. [6]. We denote by $\mathcal{M}$ the set of all random variables $\xi\colon\Omega\to\mathbb{R}$, where two of them are identified if they have the same law. Further, let $\mathcal{M}_{\infty}$ and $\mathcal{M}_{1}$ be the sets of all random variables in $\mathcal{M}$ which are essentially bounded, and have finite first moments, respectively. In the following, we analyze the main results of the previous section for several stochastic orders. For definitions and further details on stochastic orders, we refer to [21, 27].